40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 20 00:59:02 2011, kiwamu, Summary, General, Weekly report 
    Reply  Mon Jun 27 10:04:29 2011, kiwamu, Summary, General, Weekly report 
       Reply  Mon Jul 4 14:27:35 2011, kiwamu, Summary, General, Weekly report 
          Reply  Mon Jul 25 11:02:19 2011, kiwamu, Summary, General, Weekly report 
             Reply  Mon Jul 25 11:04:22 2011, kiwamu, Summary, General, Weekly report 
                Reply  Mon Jul 25 11:06:38 2011, kiwamu, Summary, General, Weekly report 
                   Reply  Mon Aug 1 22:30:59 2011, kiwamu, Summary, General, Weekly report 
Message ID: 5027     Entry time: Mon Jul 25 11:04:22 2011     In reply to: 5026     Reply to this: 5028
Author: kiwamu 
Type: Summary 
Category: General 
Subject: Weekly report 

Summary of the week ending July 17th.  Number of elog entries = 20

- LSC
 * BO switching logic for the WF was installed on c1lsc
 * Channel mapping updated

- SUS
 * Oplev health check. Spectrum of each quadrant on every suspension was inspected and looked healthy.

- OAF
 * BNS interface board was attached to an AA board
 * The AA board was installed on 1X7. The Electro-optic fanout chassis on 1X7 is now sitting on a jack, this should be fixed.

- Fiber experiment (ALS)
 * Fibre from the ETMX and EMTY tables were routed to the PSL table

- Misc.
 * Alberto came over to the 40m with Wagonga
 
ELOG V3.1.3-